OpenStudy (anonymous):

A salesman had to drive from his home to meet a client. If he drove 1/4 of the distance on Monday and 1/3 of the distance on Tuesday, what portion of the trip is left for him to travel on Wednesday?

OpenStudy (anonymous):

1 - (1/4) - (1/3) = x (12/12) - (3/12) - (4/12) = 5/12